OpenStudy (anonymous):

|2x + 3| = x + 1 can anyone explain to me why it has no solution?

OpenStudy (anonymous):

\(|2x+3|=x+1\therefore-2x-3=x+1,2x+3=x+1\) And the solutions for \(x\) do exist, unless I'm mistaken.

OpenStudy (anonymous):

@badreferences It appears you may be mistaken. http://www.wolframalpha.com/input/?i=abs%282x%2B3%29+%3D+x%2B1

OpenStudy (anonymous):

for values of 2x+3<0, -2x-3 = x+1 for values of 2x+3>=0, 2x+3 = x+1 So we solve for 2x+3<0 (x<-3/2) -2x-3=x+1 -4=3x x=-4/3 Solve for 2x+3>=0 (x>=-3/2) 2x+3 = x+1 x=-2 So the two solutions are (x=-4/3 and x=-2) But we have a problem... x = -4/3 is only valid for values of x<-3/2 -4/3 is greater than -3/2 so we cannot count this as a solution Same with x = -2. It is only valid for values of x>=-3/2 -2 is less than -3/2 so we cannot count this as a solution. Therefore there are no solutions

OpenStudy (anonymous):

It's true, I forgot to bound the values for the absolute values. My badd.

OpenStudy (anonymous):

Perhaps, you should acquire some goodreferences? :-P

OpenStudy (anonymous):

My name shall forevermore haunt me. XD

OpenStudy (anonymous):

x is -2

OpenStudy (anonymous):

remember the domain for the absolute value (refer to my post above)

OpenStudy (anonymous):

I don't understand!

OpenStudy (anonymous):

What a terrible mistake to make.

OpenStudy (anonymous):

@leemichaeljh when we are solving for non-negative values of 2x+3 (the bit inside the abs), we must take into account that if 2x+3 >= 0, then x >= -3/2. After simplifying 2x+3 = x+1, we can deduce that x = -2, but this conflicts with x >= -3/2 so this is not a solution for x.

OpenStudy (anonymous):

thx guys

OpenStudy (anonymous):

find the value of X